If 23.32g CaCO3 and 65.3 mL 6.00 M HCl are mixed, how many grams of CO2 will be formed?
1
Expert's answer
2012-10-08T09:57:21-0400
n of CaCO3=23.32/100=0.2332 mol

n of HCl=0.3918 mol
& 6 mol --------1000 ml
& x mol---------65.3 ml
x=0.3918

As can we see from reaction CaCO3 is in excess

& 0.3918 m
CaCO3 + & 2 HCl =CaCl2 + CO2 + H2O
2 44

m= 44*0.3918/2=8.62 g